# Optimal Transmission Power Scheduling for Networked Control System under DoS Attack

### v1

@@ III-B Optimal power scheduling
We omit the remaining terms of (17) and the last term of (18) as they are independent of $\pi$. Note that $e_{k}$ can be fully recovered by the local power scheduler as $\delta_{k-1}\in\mathcal{I}_{k}^{p}$. Let us apply the dynamic programming approach to (19). Synthesizing the optimal scheduler boils down to characterize the value function $V_{k}(\mathcal{I}_{0}^{p})$, which is defined as in (15). Moreover, we have
@@
Theorem 1 characterizes the optimal power scheduler for the finite-horizon problem (4). Note that it is difficult to compute $\iota_{k}$ in (13), in particular for a long horizon $T$. Alternatively, there is a practical greedy way to approximate $q^{*}_{k}$ as follows
$q_{k}^{\ast}$ $=$ $\mathop{\operatorname{argmin}}\limits_{q_{k}\in\mathcal{Q}}g(e_{k},a_{k},q_{k})$ (21)
with $g(e_{k},a_{k},q_{k})$ defined in (14). The greedy scheduling policy solved from (21) is suboptimal yet has a rather low computation complexity.
Theorem 1 assumes that the attack energy $a_{k}$ is known, the following proposition considers the case when attack energy $a_{k}$ is unknown.
Proposition 1
@@
Next, we provide an approximation to the optimal scheduler when the attack energy $a_{k}$ is unknown. It follows from the proof of Theorem 1. Then, we consider the optimisation problem (19). Taking the expectation of $p_{k}$ over the distribution $\mathcal{D}_{a}$ yields
$\mathbf{E}_{a_{k}\sim\mathcal{D}_{a}}[p(q_{k},a_{k})]=\int_{a_{k}\in\mathcal{S}}\Big{(}Q_{f}^{-1}\Big{(}\frac{q_{k}}{2}\Big{)}\Big{)}^{2}\frac{a_{k}+\sigma^{2}}{\alpha}{\rm d}a_{k}.$ (22)
Then we replace $p_{k}$ with $\mathbf{E}_{a_{k}\sim\mathcal{D}_{a}}[p(q_{k},a_{k})]$ in scheduling policy (13) and (14). Then, an approximation-based greedy scheduling is
@@ IV Infinite-horizon case
The minimum achievable packet dropout probability under the attack energy $a$, i.e., $q_{m}(a)=2Q_{f}\bigg{(}\sqrt{\frac{\alpha p_{\max}}{a+\sigma^{2}}}\bigg{)}$ satisfies $\mathbf{E}_{a\sim\mathcal{D}_{a}}[q_{m}(a)]<\frac{1}{\rho(A)^{2}}.$
@@
$\bar{\Psi}(f^{\ast},\pi)=\mathbf{E}\big{[}\sum_{k=0}^{\infty}\gamma^{k+1}w_{k}^{\top}Pw_{k}+\gamma^{k}e_{k}^{\top}\Sigma e_{k}+\lambda\gamma^{k}p_{k}\big{]}.$ (26)
@@
with the per-stage cost $g$ defined in (14), $\Sigma=L^{\top}(R+\gamma B^{\top}PB)L$ and initial value $(e_{0},a_{0}):=(e,a)$. We next formulate the optimization problem (27) as a MDP with the control model $(\mathcal{X},\mathcal{Q},\mathcal{P},g)$, where $\mathcal{X}:=\mathbb{R}^{n}\times\mathcal{S}$ denotes the state space, $\mathcal{Q}$ denotes the action space, $\mathcal{P}$ denotes the Borel measurable transition kernel defined on $(\mathcal{X},\mathcal{Q})$. Note that $w_{k}$ obeys a distribution of $\mathcal{N}(\mathbf{0},W)$. Denote $e^{+},a^{+}$ as the next states of $e$ and $a$, respectively. Denote $\mathcal{P}$ as the transition probability from $(e,a,q)$ to $(e^{+},a^{+})$. By (10), we have
@@
where $\mu_{w}$ is the probability density function of the random variable $w\sim\mathcal{N}(\mathbf{0},W)$. Note that (28) holds as the process noise $w$ is independent of attack energy $a\in\mathcal{D}_{a}$ and $q$ is the current decision.
According to the optimization problem (27), define an associated $n$-stage cost under the policy $\pi\in\Pi$:
$J_{n}(e,a,\pi):=\mathbf{E}^{\pi}\big{[}\sum_{t=0}^{n-1}\gamma^{t}g(e_{t},a_{t},q_{t})\big{]}$ (29)
with the initial value $(e,a)\in\mathcal{X}$ and $n\geq 1$. The decision variables $q_{t}$ for $t\in\mathbb{N}_{[0,n-1]}$ is chosen according to policy $\pi$. Let $G(e,a)$ be a class of nonnegative and lower semi-continuous (l.s.c.) functions on $(e,a)\in\mathcal{X}$. Define the value iteration sequence $V_{n}(e,a)\in G(e,a)$, for $n\geq 1$ and $V_{0}(\cdot,\cdot)=0$:
$V_{n}(e,a):$ $=$ $\min_{q\in\mathcal{Q}(e,a)}\big{\{}g(e,a,q)$ (30)
$\hskip 10.00002pt+\gamma\mathbf{E}[V_{n-1}(e^{+},a^{+}|e,a,q)]\big{\}},$
where $\mathcal{Q}(e,a)$ denotes a non-empty set associated to $(e,a)$, for any $(e,a)\in\mathcal{X}$. Then, we have $V_{n}(e,a)=\inf_{\pi}J_{n}(e,a,\pi)$ given initial value $(e,a)\in\mathcal{X}$, for $n\geq 1$.
@@
Let Assumption 2 hold, we have the following claims:
@@
    $V^{\ast}(e,a)=\min_{q\in\mathcal{Q}(e,a)}\Big{[}g(e,a,q)+\gamma\mathbf{E}[V^{\ast}(e^{+},a^{+})|e,a,q]\Big{]}.$ (31)
@@
    the multifunction $\pi(e,a)\rightarrow\mathcal{Q}(e,a)$ is l.s.c.;
@@
    there exists a policy $\bar{\pi}$ such that $J_{\infty}(e,a,\bar{\pi}):=\sum_{k=0}^{\infty}\mathbf{E}^{\bar{\pi}}\Big{[}\gamma^{k}g(e_{k},a_{k},\bar{q})\Big{]}<\infty$ for each $(e,a)\in\mathcal{X}$.
The first condition holds by the definition of $g(e,a,q)$, as in (14). It is inf-compact on $\mathcal{X}\times\mathcal{Q}$ as the set $\{q\in\mathcal{Q}(e,a)|g(e,a,q)\leq r\}$ is compact. The second condition holds as $\mathcal{P}(e,a,q)$ is continuous on $(e,a,q)\in\mathcal{X}\times\mathcal{A}$. Then for any continuous and bounded function $V^{\ast}(e,a)$ on $(e,a,q)$, the map $(e,a,q)\rightarrow\inf_{\mathcal{X}}J^{\ast}(e,a)\mathcal{P}(e^{+},a^{+}|e,a,q){\rm d}e^{+}{\rm d}a^{+}$ is continuous on $(e,a,q)\in\mathcal{X}\times\mathcal{A}$. The third condition holds as $\mathcal{X}\times\mathcal{Q}$ is convex, see [23].
To verify the last condition, we choose $\bar{\pi}(\mathcal{I}_{k}^{p})=p_{\max}$ for all $k$. Then, $q(a_{k},p_{\max})=2Q_{f}\Big{(}\sqrt{\frac{\alpha p_{\max}}{a_{k}+\sigma^{2}}}\Big{)}$. The expected dropout probability under $\bar{\pi}$ is
$\bar{q}=\mathbf{E}_{a\sim\mathcal{D}_{a}}[q(a,p_{\max})].$ (32)
The next is to calculate $J_{\infty}(e,a,\bar{\pi})$. Denote $X_{k}^{e}:=\mathbf{E}[e_{k}^{\top}e_{k}]$. By the expression of estimation error as in (10), we have
$X_{k}^{e}=\left\{\begin{array}[]{ll}{\rm{tr}}(X_{0})&{\rm if}~{}k=0\\ {\rm{tr}}(W)&{\rm if}~{}k>0~{}\&~{}\delta_{k-1}=1\\ A^{\top}X_{k-1}^{e}A+W&{\rm if}~{}k>0~{}\&~{}\delta_{k-1}=0.\end{array}\right.$ (36)
Accordingly, define $\theta_{t}^{w}:={\rm{tr}}\big{(}\sum_{r=0}^{t-1}(A^{r})^{\top}\Sigma A^{r}W\big{)}$, $\theta_{t}^{x}:={\rm{tr}}\big{(}(A^{t})^{\top}\Sigma A^{t}X_{0}\big{)}+\theta_{t}^{w},\quad{\rm for}~{}t\geq 1.$ When $k=0$, $\mathbf{E}^{\bar{\pi}}[\gamma^{k}e_{k}^{\top}\Sigma e_{k}]={\rm{tr}}(\Sigma X_{0})$. When $k=1$,
$\mathbf{E}^{\bar{\pi}}[\gamma^{k}e_{k}^{\top}\Sigma e_{k}]$ $=$ $\gamma(\bar{q}\theta_{1}^{x}+(1-\bar{q})\theta_{1}^{w})$
$=$ $\gamma{\rm{tr}}(\Sigma W+\bar{q}A^{\top}\Sigma AX_{0}).$
When $k=2$,
$\mathbf{E}^{\bar{\pi}}[\gamma^{k}e_{k}^{\top}\Sigma e_{k}]$ $=$ $\gamma^{2}\big{(}(1-\bar{q})\theta_{1}^{w}+\bar{q}(1-\bar{q})\theta_{2}^{w}+\bar{q}^{2}\theta_{2}^{x}\big{)}$
$=$ $\gamma^{2}{\rm{tr}}(\Sigma W+\bar{q}A^{\top}\Sigma AW+\bar{q}^{2}(A^{2})^{\top}\Sigma A^{2}X_{0}).$
By induction, for all $k\geq 1$, we have
$\mathbf{E}^{\bar{\pi}}[\gamma^{k}e_{k}^{\top}\Sigma e_{k}]$ (37)
$=$ $\gamma^{k}{\rm{tr}}\Big{(}\sum_{r=0}^{k-1}\bar{q}^{k}(A^{k})^{\top}\Sigma A^{k}W+\bar{q}^{k}(A^{k})^{\top}\Sigma A^{k}X_{0}\Big{)}.$
@@
$\mathbf{E}^{\bar{\pi}}\bigg{[}\sum_{k=0}^{\infty}\gamma^{k}e_{k}^{\top}\Sigma e_{k}\bigg{]}=\frac{\gamma}{1-\gamma}{\rm{tr}}\bigg{(}\sum_{r=0}^{\infty}(\gamma\bar{q})^{r}(A^{r})^{\top}\Sigma A^{r}W\bigg{)}$
$\hskip 90.00014pt+{\rm{tr}}\bigg{(}\sum_{r=0}^{\infty}(\gamma\bar{q})^{r}(A^{r})^{\top}\Sigma A^{r}X_{0}\bigg{)}.$ (38)
Note that when $\gamma\bar{q}(\rho(A))^{2}<1$, $\sum_{k=0}^{\infty}\mathbf{E}^{\bar{\pi}}[\gamma^{k}e_{k}^{\top}\Sigma e_{k}]<\infty$. Then Assumption 2 is sufficient for its establishment. Moreover, we have
$\mathbf{E}^{\bar{\pi}}\bigg{[}\sum_{k=0}^{\infty}\gamma^{k}e_{k}^{\top}\Sigma e_{k}\bigg{]}-\mathbf{E}^{\bar{\pi}}\bigg{[}\sum_{k=0}^{\infty}\gamma^{k+1}q_{k}e_{k}^{\top}A^{\top}\Sigma Ae_{k}\bigg{]}$ (39)
$=$ ${\rm{tr}}(\Sigma X_{0})+\frac{\gamma}{1-\gamma}{\rm{tr}}(\Sigma W).$
The total transmission cost under power scheduling law $\bar{\pi}$ is written as
$J_{p}^{\bar{\pi}}=\sum_{k=0}^{\infty}\gamma^{k}p_{\max}=\frac{p_{\max}}{1-\gamma}.$ (40)
Combine (IV), (39) and (40), we obtain that
$J_{\infty}(e,a,\bar{\pi})=\mathbf{E}^{\bar{\pi}}\bigg{[}\sum_{k=0}^{\infty}\gamma^{k+1}q_{k}e_{k}^{\top}A^{\top}\Sigma Ae_{k}\bigg{]}+\lambda J_{p}^{\bar{\pi}}<\infty.$
Conditions 1-4) are verified. According to [22], claims 1-3) holds. $\blacksquare$
The following theorem develops the optimal scheduler and the cost achieved by the designed scheduling policy.
@@
Consider the optimization problem (6) for system (1). Fix the optimal control law as the certainty equivalence controller (25). Let Assumption 2 hold and let
$q^{\ast}(e,a)=\mathop{\operatorname{argmin}}_{q\in\mathcal{Q}}\big{\{}g(e,a,q)+\gamma\mathbf{E}[V^{\ast}(e^{+},a^{+})|e,a,q]\big{\}},$ (41)
where $V^{\ast}$ is solved from (31).
@@
    $p=\pi^{*}(q^{\ast}(e,a),a)=\Big{(}Q_{f}^{-1}\Big{(}\frac{q^{\ast}(e,a)}{2}\Big{)}\Big{)}^{2}\frac{a+\sigma^{2}}{\alpha}.$ (42)
@@
    Let $\Theta$ be the solution of Laypunov equation:
    $\gamma\bar{q}A^{\top}\Theta A+\Sigma=\Theta.$ (43)
    where $\bar{q}$ is given by (32). The total cost achieved by the optimal power scheduler and the optimal control is upper-bounded by
    $\bar{\Psi}(f^{\ast},\pi^{\ast})$$\leq$ $\frac{\gamma{\rm{tr}}(PW+\Theta W)+\lambda p_{\max}}{1-\gamma}+{\rm{tr}}(\Theta X_{0}).$ (44)
Proof. The first claim follows from Proposition 2. By Assumption 2, the solution of (43) is written as
$\Theta=\sum_{k=0}^{\infty}(\gamma\bar{q})^{k}(A^{k})^{\top}\Sigma A^{k}.$ (45)
Substituting (45) into (IV), we have
$\mathbf{E}\bigg{[}\sum_{k=0}^{\infty}\gamma^{k}e_{k}^{\top}\Sigma e_{k}\bigg{]}$ $=$ $\frac{\gamma}{1-\gamma}{\rm{tr}}(\Theta W)+{\rm{tr}}(\Theta X_{0}).$ (46)
Substitute (46) and (40) into (26), we have
$\bar{\Psi}(f^{\ast},\pi^{*})$$\leq$ $\mathbf{E}\bigg{[}\sum_{k=0}^{\infty}\gamma^{k+1}w_{k}^{\top}Pw_{k}\bigg{]}$
$+\frac{\gamma{\rm{tr}}(\Theta W)}{1-\gamma}+{\rm{tr}}(\Theta X_{0})+\frac{\lambda p_{\max}}{1-\gamma}$
$=$ $\frac{\gamma{\rm{tr}}(PW+\Theta W)+\lambda p_{\max}}{1-\gamma}+{\rm{tr}}(\Theta X_{0}).$
Thus, we obtain the second conclusion. $\blacksquare$
The optimization problem 2 considers discounted criteria, the minimum expected cost (44) is affected by the attack energy distribution $\mathcal{D}_{a}$, the discount factor $\gamma$, the covariance of state initial value $X_{0}$. This aligns with the result of the discounted optimization problem, see [22].

### v2

@@ III-B Optimal power scheduling
where $g(e_{k},a_{k},q_{k})$ is defined in (13). We omit the remaining terms of (16) and the last term of (17) as they are independent of $\pi$. Note that $e_{k}$ can be fully recovered by the local power scheduler as $\delta_{k-1}\in\mathcal{I}_{k}^{p}$. Let us apply the dynamic programming approach to (18). Synthesizing the optimal scheduler boils down to solving the value function $V_{0}(\mathcal{I}_{0}^{p})$, as defined in (14). Moreover,
@@
Theorem 1 characterizes the optimal power scheduler for the finite-horizon problem (4). Note that it is difficult to compute $\iota_{k}$ in (12), in particular for a long horizon $T$. An alternative greedy way to approximate $q^{*}_{k}$ is letting
$q_{k}^{\ast}$ $=$ $\mathop{\operatorname{argmin}}\limits_{q_{k}\in\mathcal{Q}}g(e_{k},a_{k},q_{k}).$ (20)
The greedy scheduling policy solved from (20) is suboptimal yet has a rather low computation complexity.
Lemma 1 and Theorem 1 assume that the attack energy $a_{k}$ is known, the following result considers the case when attack energy $a_{k}$ is unknown.
Lemma 2
@@
Next, we provide an approximation to the optimal scheduler when the attack energy $a_{k}$ is unknown. It follows from the proof of Theorem 1. Then, we consider the optimization problem (18). Taking the expectation of $p_{k}$ over the distribution $\mathcal{D}_{a}$ yields $\mathbf{E}_{a_{k}\sim\mathcal{D}_{a}}[p(q_{k},a_{k})]$. Then we replace $p_{k}$ with $\mathbf{E}_{a_{k}\sim\mathcal{D}_{a}}[p(q_{k},a_{k})]$ in scheduling policy (12) and (13). Then, an approximation-based greedy scheduling is
@@
Remark 2
The greedy policy (20) requires real-time knowledge of the attack energy $a_{k}$ and does not rely on any distributional information about the attack energy, i.e., conditions 1) and 3) in Assumption 1, whereas (21) works the other way around. Furthermore, since the $p(q_{k},a_{k})$ given in (13) is linear in $a_{k}$, the cost function (18) achieved by the greedy scheduler (20) remains consistent across various attack energy distributions with the same mean. Additionally, since the average of greedy strategy (20) with respect to the random variable $a_{k}$ is identical to the approximation-based greedy strategy (21), the performance of strategy (21) converges to that of strategy (20) as the time horizon $T$ increases. This finding will be demonstrated in the Simulation section.
@@ IV Infinite-horizon case
The minimum achievable packet dropout probability under the attack energy $a$, i.e., $q_{m}(a)=2Q_{f}\bigg{(}\sqrt{\frac{\alpha p_{\max}}{a+\sigma^{2}}}\bigg{)}$, satisfies $\mathbf{E}_{a\sim\mathcal{D}_{a}}[q_{m}(a)]<\frac{1}{\rho(A)^{2}}.$
@@
$\hat{\Psi}(f^{\ast},\pi)=\mathbf{E}\big{[}\sum_{k=0}^{\infty}\gamma^{k+1}w_{k}^{\top}Pw_{k}+\gamma^{k}e_{k}^{\top}\Sigma e_{k}+\lambda\gamma^{k}p_{k}\big{]}.$ (24)
@@
with the per-stage cost $g$ defined in (13), $\Sigma=L^{\top}(R+\gamma B^{\top}PB)L$ and initial value $(e_{0},a_{0}):=(e,a)$. We next formulate the optimization problem (25) as a MDP with the control model $(\mathcal{X},\mathcal{Q},\mathcal{P},g)$, where $\mathcal{X}:=\mathbb{R}^{n}\times\mathcal{S}$ denotes the state space, $\mathcal{Q}$ denotes the action space, $\mathcal{P}$ denotes the Borel measurable transition kernel defined on $(\mathcal{X},\mathcal{Q})$. Note that $w_{k}$ obeys a distribution of $\mathcal{N}(\mathbf{0},W)$. Denote $e^{+},a^{+}$ as the next states of $e$ and $a$, respectively. Then, by (10), the transition probability from $(e,a,q)$ to $(e^{+},a^{+})$ is given by
@@
where $\mu_{w}$ is the p.d.f. of the random variable $w\sim\mathcal{N}(\mathbf{0},W)$. Note that (26) holds as the process noise $w$ is independent of attack energy $a\in\mathcal{D}_{a}$ and $q$ is the current decision.
According to the optimization problem (25), define an associated $n$-stage cost under the policy $\pi\in\Pi$: $J_{n}(e,a,\pi):=\mathbf{E}^{\pi}\big{[}\sum_{t=0}^{n-1}\gamma^{t}g(e_{t},a_{t},q_{t})\big{]}$ with the initial value $(e,a)\in\mathcal{X}$ and $n\geq 1$. The decision variables $q_{t}$ for $t\in\mathbb{N}_{[0,n-1]}$ is chosen according to the policy $\pi$. Let $G(e,a)$ be a class of nonnegative and lower semi-continuous (l.s.c.) functions on $(e,a)\in\mathcal{X}$. Define the value iteration sequence $V_{n}(e,a)\in G(e,a)$, for $n\geq 1$ and $V_{0}(\cdot,\cdot)=0$:
$V_{n}(e,a)$ $=$ $\min_{q\in\mathcal{W}(e,a)}\big{\{}g(e,a,q)+\gamma\mathbf{E}[V_{n-1}(e^{+},a^{+}|e,a,q)]\big{\}},$
where $\mathcal{W}(e,a)$ highlights that, for any $(e,a)\in\mathcal{X}$, a non-empty set is associated to $(e,a)$. Then, we have $V_{n}(e,a)=\inf_{\pi}J_{n}(e,a,\pi)$ given the initial value $(e,a)\in\mathcal{X}$, for $n\geq 1$.
@@
Let Assumptions 1 and 2 hold, we have the following claims:
@@
    $V^{\ast}(e,a)=\min_{q\in\mathcal{W}(e,a)}\Big{[}g(e,a,q)+\gamma\mathbf{E}[V^{\ast}(e^{+},a^{+})|e,a,q]\Big{]}.$ (27)
@@
    the multifunction $(e,a)\rightarrow\mathcal{W}(e,a)$ is l.s.c.;
@@
    there exists a policy $\hat{\pi}$ such that $J_{\infty}(e,a,\hat{\pi}):=\sum_{k=0}^{\infty}\mathbf{E}^{\hat{\pi}}\Big{[}\gamma^{k}g(e_{k},a_{k},\hat{q})\Big{]}<\infty$ for each $(e,a)\in\mathcal{X}$, where $\hat{q}=\hat{\pi}(\mathcal{I}_{k}^{p})$.
The first condition holds by the definition of $g(e,a,q)$, as in (13). It is inf-compact on $\mathcal{X}\times\mathcal{Q}$ as the set $\{q\in\mathcal{W}(e,a)|g(e,a,q)\leq r\}$ is compact. The second condition holds as $\mathcal{P}(e,a,q)$ is continuous on $(e,a,q)\in\mathcal{X}\times\mathcal{A}$. Condition 2) implies that, for any continuous and bounded function $V^{\ast}(e,a)$ on $(e,a,q)$, the map $(e,a,q)\rightarrow\int_{\mathcal{X}}V^{\ast}(e,a)\mathcal{P}(e^{+},a^{+}|e,a,q){\rm d}e^{+}{\rm d}a^{+}$ is continuous on $(e,a,q)\in\mathcal{X}\times\mathcal{A}$. The third condition holds as $\mathcal{X}\times\mathcal{Q}$ is convex, see [35].
To verify the last condition, we choose $\hat{\pi}(\mathcal{I}_{k}^{p})=p_{\max}$ for all $k$. Then, the expected dropout probability under $\hat{\pi}$ is $\hat{q}=\mathbf{E}_{a\sim\mathcal{D}_{a}}[q(a,p_{\max})]$ with $q(a,p_{\max})=2Q_{f}\Big{(}\sqrt{\frac{\alpha p_{\max}}{a+\sigma^{2}}}\Big{)}$. The next is to calculate $J_{\infty}(e,a,\hat{\pi})$. Define $\theta_{t}^{w}:={\rm{tr}}\big{(}\sum_{r=0}^{t-1}(A^{r})^{\top}\Sigma A^{r}W\big{)}$, $\theta_{t}^{x}:={\rm{tr}}\big{(}(A^{t})^{\top}\Sigma A^{t}X_{0}\big{)}+\theta_{t}^{w},\quad{\rm for}~{}t\geq 1.$ When $k=0$, $\mathbf{E}^{\hat{\pi}}[\gamma^{k}e_{k}^{\top}\Sigma e_{k}]={\rm{tr}}(\Sigma X_{0})$. When $k=1$, $\mathbf{E}^{\hat{\pi}}[\gamma^{k}e_{k}^{\top}\Sigma e_{k}]=\gamma(\hat{q}\theta_{1}^{x}+(1-\hat{q})\theta_{1}^{w})=\gamma{\rm{tr}}(\Sigma W+\hat{q}A^{\top}\Sigma AX_{0}).$ When $k=2$, $\mathbf{E}^{\hat{\pi}}[\gamma^{k}e_{k}^{\top}\Sigma e_{k}]=\gamma^{2}\big{(}(1-\hat{q})\theta_{1}^{w}+\hat{q}(1-\hat{q})\theta_{2}^{w}+\hat{q}^{2}\theta_{2}^{x}\big{)}=\gamma^{2}{\rm{tr}}(\Sigma W+\hat{q}A^{\top}\Sigma AW+\hat{q}^{2}(A^{2})^{\top}\Sigma A^{2}X_{0}).$ By induction, for all $k\geq 1$, we have
$\mathbf{E}^{\hat{\pi}}[\gamma^{k}e_{k}^{\top}\Sigma e_{k}]$ (28)
$=$ $\gamma^{k}{\rm{tr}}\Big{(}\hat{q}^{k}(A^{k})^{\top}\Sigma A^{k}X_{0}+\sum_{r=0}^{k-1}\hat{q}^{r}(A^{r})^{\top}\Sigma A^{r}W\Big{)}.$
@@
$\mathbf{E}^{\hat{\pi}}\bigg{[}\sum_{k=0}^{\infty}\gamma^{k}e_{k}^{\top}\Sigma e_{k}\bigg{]}=\frac{\gamma}{1-\gamma}{\rm{tr}}\bigg{(}\sum_{r=0}^{\infty}(\gamma\hat{q})^{r}(A^{r})^{\top}\Sigma A^{r}W\bigg{)}$
$\hskip 90.00014pt+{\rm{tr}}\bigg{(}\sum_{r=0}^{\infty}(\gamma\hat{q})^{r}(A^{r})^{\top}\Sigma A^{r}X_{0}\bigg{)}.$ (29)
When $\gamma\hat{q}(\rho(A))^{2}<1$, we have $\sum_{k=0}^{\infty}\mathbf{E}^{\hat{\pi}}[\gamma^{k}e_{k}^{\top}\Sigma e_{k}]<\infty$. Then Assumption 2 is sufficient for its establishment. Moreover, we have
$\mathbf{E}^{\hat{\pi}}\bigg{[}\sum_{k=0}^{\infty}\gamma^{k}e_{k}^{\top}\Sigma e_{k}\bigg{]}-\mathbf{E}^{\hat{\pi}}\bigg{[}\sum_{k=0}^{\infty}\gamma^{k+1}\hat{q}_{k}e_{k}^{\top}A^{\top}\Sigma Ae_{k}\bigg{]}$ (30)
$=$ ${\rm{tr}}(\Sigma X_{0})+\frac{\gamma}{1-\gamma}{\rm{tr}}(\Sigma W),$
which holds by substituting (10) into (30). The total transmission cost under the power scheduling law $\hat{\pi}$ is
$J_{p}^{\hat{\pi}}=\sum_{k=0}^{\infty}\gamma^{k}p_{\max}=\frac{p_{\max}}{1-\gamma}.$ (31)
Combining (IV), (30) and (31), we obtain that
$J_{\infty}(e,a,\hat{\pi})=\mathbf{E}^{\hat{\pi}}\bigg{[}\sum_{k=0}^{\infty}\gamma^{k+1}\hat{q}_{k}e_{k}^{\top}A^{\top}\Sigma Ae_{k}\bigg{]}+\lambda J_{p}^{\hat{\pi}}<\infty.$
According to [34], claims 1-3) holds. $\blacksquare$
The following theorem develops the optimal scheduler and analyzes its performance.
@@
Consider the optimization problem (6) for system (1). Fix the optimal control law as the certainty equivalence controller (23). Let Assumptions 1 and 2 hold and let
$q^{\ast}(e,a)=\mathop{\operatorname{argmin}}_{q\in\mathcal{Q}}\big{\{}g(e,a,q)+\gamma\mathbf{E}[V^{\ast}(e^{+},a^{+})|e,a,q]\big{\}},$
where $V^{\ast}(e,a)$ is solved from Bellman equation (27).
@@
    $p=\pi^{\ast}(e,a)=\Big{(}Q_{f}^{-1}\Big{(}\frac{q^{\ast}(e,a)}{2}\Big{)}\Big{)}^{2}\Big{(}\frac{a+\sigma^{2}}{\alpha}\Big{)}.$ (32)
@@
    Let $\tilde{q}$ be the solution of
    $\operatorname{argmin}_{q\in\mathcal{Q}}\bigg{\{}\frac{\gamma{\rm{tr}}(\Theta W)}{1-\gamma}+{\rm{tr}}(\Theta X_{0})+\frac{\lambda\mathbf{E}_{a\sim\mathcal{D}_{a}}[p(q,a)]}{1-\gamma}\bigg{\}},$
    ${\mathrm{s.t.}}~{}\gamma qA^{\top}\Theta A+\Sigma=\Theta.$ (33)
    The total cost achieved by the optimal power scheduler and the optimal control is upper-bounded by
    $\hat{\Psi}(f^{\ast},\pi^{\ast})\leq\frac{\gamma{\rm{tr}}(PW+\tilde{\Theta}W)+\lambda\tilde{p}}{1-\gamma}+{\rm{tr}}(\tilde{\Theta}X_{0}),$ (34)
    where $\tilde{p}=\mathop{\mathbf{E}}\limits_{a_{k}\sim\mathcal{D}_{a}}[p(\tilde{q},a_{k})]$ and $\tilde{\Theta}$ satisfies
    $\gamma\tilde{q}A^{\top}\tilde{\Theta}A+\Sigma=\tilde{\Theta}.$ (35)
Proof. The first claim follows from Proposition 1. Next, we show the second claim. Define a class of constant-power scheduling law $\bar{\Pi}$ and denote $\tilde{\pi}\in\bar{\Pi}$ as the optimal constant-power scheduling law minimizing the total cost (24). Given Assumption 2,
$\tilde{\Theta}=\sum_{k=0}^{\infty}(\gamma\tilde{q})^{k}(A^{k})^{\top}\Sigma A^{k}.$ (36)
is the solution to Lyapunov equation (35). Substituting (36) into (IV), we have
$\mathbf{E}^{\tilde{\pi}}\bigg{[}\sum_{k=0}^{\infty}\gamma^{k}e_{k}^{\top}\Sigma e_{k}\bigg{]}$ $=$ $\frac{\gamma}{1-\gamma}{\rm{tr}}(\tilde{\Theta}W)+{\rm{tr}}(\tilde{\Theta}X_{0}).$ (37)
Substituting (37) and $J_{p}^{\tilde{\pi}}=\sum_{k=0}^{\infty}\gamma^{k}\tilde{p}=\frac{\tilde{p}}{1-\gamma}$ into (24), we have
$\hskip 10.00002pt\hat{\Psi}(f^{\ast},\tilde{\pi})$
$=\mathbf{E}^{\tilde{\pi}}\bigg{[}\sum_{k=0}^{\infty}\gamma^{k+1}w_{k}^{\top}Pw_{k}\bigg{]}+\frac{\gamma{\rm{tr}}(\tilde{\Theta}W)}{1-\gamma}+{\rm{tr}}(\tilde{\Theta}X_{0})+\frac{\lambda\tilde{p}}{1-\gamma}$
$=\frac{\gamma{\rm{tr}}(PW+\tilde{\Theta}W)+\lambda\tilde{p}}{1-\gamma}+{\rm{tr}}(\tilde{\Theta}X_{0}).$
Thus, $\tilde{p}=\tilde{\pi}(e,a)$ solved from (2) is the optimal constant power, where $\frac{\gamma{\mathrm{tr}}(PW)}{1-\gamma}$ is omitted as it is independent of scheduling policy. Note that the scheduling policy (32) is optimal among the class of policies depending on primitive random variables described by $\mathcal{I}_{k}^{p}$, including $\bar{\Pi}$. As a consequence, the performance achieved by the scheduling policy $\pi^{\ast}$ is upperbounded by that achieved by $\tilde{\pi}$, i.e., $\hat{\Psi}(f^{\ast},\pi^{\ast})\leq\hat{\Psi}(f^{\ast},\tilde{\pi})$. Thus, we obtain the second conclusion. $\blacksquare$
Since Problem 2 considers discounted criteria, the minimum expected cost (34) is affected by the attack energy distribution $\mathcal{D}_{a}$, the discount factor $\gamma$, the covariance of state initial value $X_{0}$. This aligns with the result of the discounted optimization problem, see [34].